### v2

# Mixing It Up: Exploring Mixer Networks for Irregular Multivariate Time Series Forecasting

Abstract
Forecasting irregularly sampled multivariate time series with missing values (IMTS) is a fundamental challenge in domains such as healthcare, climate science, and biology. While recent advances in vision and time series forecasting have shown that lightweight MLP-based architectures (e.g., MLP-Mixer, TSMixer) can rival attention-based models in both accuracy and efficiency, their applicability to irregular and sparse time series remains unexplored. In this paper, we propose IMTS-Mixer, a novel architecture that adapts the principles of Mixer models to the IMTS setting. IMTS-Mixer introduces two key components: (1) ISCAM, a channel-wise encoder that transforms irregular observations into fixed-size vectors using simple MLPs, and (2) ConTP, a continuous time decoder that supports forecasting at arbitrary time points. We evaluate IMTS-Mixer on two challenging IMTS benchmarks—PMAU and Physiome-ODE—covering a total of 54 datasets. Our model achieves state-of-the-art performance in both forecasting accuracy and inference time, while using significantly fewer parameters than existing baselines.
Codehttps://anonymous.4open.science/r/IMTS-Mixer-682E
1 Introduction
Time series forecasting is a critical task across various domains. While a significant body of research has focused on forecasting regularly sampled, fully observed multivariate time series, many real-world applications, such as healthcare, economics, astronomy, and climate, deal with time series data that are irregularly sampled, with observations missing across different channels and timestamps. When such data are aligned to a global timeline, the result is a highly sparse and uneven multivariate structure, as illustrated in Figure 1. We refer to such data as irregularly sampled multivariate time series with missing values (IMTS).
[…]
Figure 1: Example of an IMTS Forecasting task. The observations and forecasting targets are irregularly spaced.
PMAU (ICML ’24)
Avg. Rank %Gap-MSE #Params. Inf. Time
GraFITi 2 2.9% 313k 2.2s
tPATCHGNN 2.75 5.2% 265k 4.2s
TimeCHEAT 9.75 31% 751k 46s
IMTS-Mixer 1.25 1.3% 204k 1.3s
Physiome-ODE (ICLR ’25)
Avg. Rank %Gap-MSE #Params. Inf. Time
GraFITi 3.20 42% 883k 0.32s
tPATCHGNN 5.94 79% 287k 0.14s
TimeCHEAT 5.86 141% 566k 2.5s
IMTS-Mixer 1.20 4.1% 136k 0.07s
Table 1: Aggregated results for selected models on two benchmarks: PMAU (Zhang et al. 2024a) and Physiome-ODE (Klötergens et al. 2024). We report the average ranks by MSE test error (Avg. Rank) and the average %-Gap to the best performing model (%Gap-MSE). Additionally, we show the average number of parameters (#Params.) and inference time for the complete test set of each sub-task (Inf. Time). For all metrics, lower is better.
Traditional approaches to modeling IMTS have relied on Neural-ODE frameworks, where the time series is modeled as the solution to an ordinary differential equation with a neural network-defined vector field (Rubanova, Chen, and Duvenaud 2019; De Brouwer et al. 2019; Schirmer et al. 2022). While powerful in principle, these models tend to be computationally expensive due to the sequential nature of ODE solvers, and they lack a systematic way to handle missing values.
Recently, attention-based models have emerged as a more efficient and accurate alternative, achieving competitive performance on IMTS forecasting tasks (Yalavarthi et al. 2024a; Zhang et al. 2024a; Liu, Cao, and Chen 2025). These models typically transform irregular channel-wise observations into fixed-size representations and apply attention mechanisms to capture inter- and intra-channel dependencies. However, such models are often resource-intensive, requiring a large number of parameters and high memory requirements.
Inspired by advances in computer vision, particularly the success of the MLP-Mixer architecture, recent work has demonstrated that carefully arranged MLP-based models can rival or even surpass attention-based models (Tolstikhin et al. 2021). This insight has also been applied to forecasting tasks with regularly sampled time series (Ekambaram et al. 2023; Chen et al. 2023). However, this line of research has not yet been extended to irregularly sampled time series with missing values.
In this paper, we introduce IMTS-Mixer, a novel architecture for forecasting with IMTS that combines the simplicity of MLP-based designs with specialized modules to address irregular sampling and forecasting at arbitrary time points. First, we propose the Irregularly Sampled Channel Aggregation Module (ISCAM), a lightweight encoder that maps each irregular channel into a fixed-size representation using simple MLPs. These embeddings are then concatenated and processed by an MLP-Mixer-style architecture to learn inter- and intra-channel interactions. Second, to support continuous-time forecasting, we introduce the Continuous Temporal Projection (ConTP) module, a decoder that takes the mixer output, a continuous-valued target timestamp, and a channel identifier to generate a prediction.
In an extensive comparison, we evaluate IMTS-Mixer on two comprehensive benchmarks:
  1. 1.
    PMAU (Zhang et al. 2024a), consisting of four IMTS datasets (Physionet 2012, MIMIC-III, Activity, and USHCN).
  2. 2.
    Physiome-ODE (Klötergens et al. 2024), a benchmark of 50 IMTS datasets derived from biological ODE models.
As shown in Table 1, IMTS-Mixer achieves state-of-the-art forecasting performance, ranking first across both benchmarks.
Our contributions are summarized as follows:
  • We propose IMTS-Mixer, a novel MLP-based architecture tailored for irregularly sampled multivariate time series with missing values (IMTS), combining efficient channel aggregation and continuous-time prediction.
  • We introduce ISCAM, a simple and effective channel-wise encoder that maps irregularly observed channels to fixed-size representations, enabling MLP processing.
  • We develop a lightweight temporal projection module (ConTP) that maps a fixed-size representation of an input channel to future IMTS predictions.
  • We conduct extensive experiments on two challenging IMTS benchmarks—PMAU and Physiome-ODE—covering a total of 54 datasets—and demonstrate that IMTS-Mixer outperforms state-of-the-art models in both forecasting accuracy and computational efficiency.
2 Related Work
2.1 Regular Multivariate Forecasting
The Regular Multivariate Time Series forecasting literature recently focused on transformer-based (Vaswani et al. 2017) architectures (Zhou et al. 2021; Wu et al. 2021; Zhou et al. 2022; Nie et al. 2022; Liu et al. 2023). However, Zeng et al. showed that a simple model based on linear layers (DLinear (Zeng et al. 2023)) can be competitive and partially outperform some of the more complex transformer architectures. Additionally, mixer architectures have been adapted from the computer vision domain (Tolstikhin et al. 2021) to time series (Ekambaram et al. 2023; Chen et al. 2023). They exclusively rely on fully-connected networks that are alternately applied along the channel and time dimension and achieve state-of-the-art results.
2.2 IMTS Forecasting
Initially, researchers relied on Neural ODE-based (Chen et al. 2018) approaches to model IMTS (Rubanova, Chen, and Duvenaud 2019; De Brouwer et al. 2019; Biloš et al. 2021; Schirmer et al. 2022). ODEs are the prevalent framework in science and engineering to model how systems evolve over time in a continuous manner. However, using neural networks to represent ODE systems has proven to be less effective than other IMTS modeling approaches and also incurs relatively long inference times (Yalavarthi et al. 2024a; Zhang et al. 2024a).
Graph Transformers for IMTS forecasting.
Apart from the ODE-based approaches the Transformable Patch Graph Neural Network (tPatchGNN) (Zhang et al. 2024a) applies a patching mechanism to each channel. The patches are then processed with a Transformer and serve as input for a Graph Neural Network (GNN) that models inter-channel dependencies. Additionally, we want to highlight Graphs for Forecasting Irregular Multivariate Time Series (GraFITi) (Yalavarthi et al. 2024a). Here, the IMTS is represented with a graph in which channels and time steps are nodes and observed values are stored in the edges between the corresponding time and channel nodes. Forecasts are made by predicting the edge values of query-time nodes with graph attention layers. While this approach provides state-of-the-art performance, it also comes with major drawbacks that relate to the graph structure itself. GraFITi relies on time nodes that are connected to multiple channel nodes, corresponding to observation steps that observe more than one channel. If only one channel is observed at a time, the resulting graph is disconnected, rendering GraFITi incapable of modeling inter-channel interactions. Another problem of GraFITi lies in the fact that the query nodes are indirectly connected and hence influence each other. This entails that predictions will change depending on what other queries we give to the model.
TimeCHEAT (Liu, Cao, and Chen 2025) transforms an IMTS into regularly observed patches using a GraFITi-based encoder. These patches are then further processed by a transformer and a linear decoder to obtain the final predictions.
3 Mixer Networks
MLP-Mixer.
The MLP-Mixer (Tolstikhin et al. 2021) was originally proposed as a simple and efficient alternative to convolutional neural networks (CNNs) and Vision Transformers (ViTs) for image classification tasks. In this architecture, a 2D image ($\in\mathbb{R}^{W\times H}$) where $W,H$ are width and height) is divided into $P$ non-overlapping patches, which are flattened and projected into vectors of length $C$, resulting in an input representation $X\in\mathbb{R}^{P\times C}$.
The core of the MLP-Mixer consists of a stack of mixer layers, each containing two separate MLP blocks:
  • a token-mixing MLP $\text{MLP}_{P}:\mathbb{R}^{P}\rightarrow\mathbb{R}^{P}$ operating across patches (rows of $X$),
  • a channel-mixing MLP $\text{MLP}_{C}:\mathbb{R}^{C}\rightarrow\mathbb{R}^{C}$ operating across channels (columns of $X$).
Each MLP block is preceded by layer normalization and followed by a residual connection. Formally, given input $X\in\mathbb{R}^{P\times C}$, the output of a single mixer layer is computed as:
$U_{*,c}$ $=X_{*,c}+\text{LayerNorm}\left(\text{MLP}_{C}(X_{*,c})\right),\;\text{ }c=1,\ldots,C$
$Z_{p,*}$ $=U_{p,*}+\text{LayerNorm}\left(\text{MLP}_{P}(U_{p,*})\right),\;\text{ }p=1,\ldots,P$
Here, $X_{*,c}$ denotes the $c$-th column (channel vector), and $X_{p,*}$ denotes the $p$-th row (patch vector).
Despite its simplicity and lack of attention mechanisms, MLP-Mixer has demonstrated competitive performance, often outperforming more complex models such as Vision Transformers (Dosovitskiy et al. 2020) on various image classification benchmarks.
TSMixer.
Motivated by the success of MLP-Mixer in vision, Chen et al. (2023) proposed TSMixer, an adaptation of the Mixer architecture for regular multivariate time series forecasting. The goal is to predict the next $Q$ time steps $Y\in\mathbb{R}^{Q\times C}$ from a past window of $T$ time steps $X\in\mathbb{R}^{T\times C}$, where $C$ denotes the number of variables (channels).
TSMixer applies a stack of Mixer layers to the input $X$, where:
  • The temporal mixer (analogous to token-mixing in images) mixes information across time steps:
    $\text{MLP}_{T}:\mathbb{R}^{T}\rightarrow\mathbb{R}^{T}$,
  • The channel mixer operates across variables:
    $\text{MLP}_{C}:\mathbb{R}^{C}\rightarrow\mathbb{R}^{C}$.
Similar to MLP-Mixer, these mixers are applied alternately with residual connections and layer normalization. After the stack of mixer layers, a final linear projection $\mathbb{R}^{T}\rightarrow\mathbb{R}^{Q}$ is applied across the temporal axis to produce the forecast $Y\in\mathbb{R}^{Q\times C}$.
4 The IMTS Forecasting Problem
Irregularly Sampled Multivariate Time Series (IMTS).
Following Shukla and Marlin (2020), we define an irregularly sampled multivariate time series (IMTS) as a collection of $C$ unaligned, sparse, and independently sampled univariate time series, one per channel:
$X\coloneqq[X_{1},\ldots,X_{C}]\in\mathcal{X}$
where each channel $X_{c}\in\textnormal{Seq}(\mathbb{R}\times\mathbb{R})$ is a sequence of observation tuples
$X_{c}\coloneqq\left(\left(t_{c,i},v_{c,i}\right)\right)_{i=1}^{N_{c}}$
with $t_{c,i}\in\mathbb{R}$: the timestamp of the $i$-th observation in channel $c$ and $v_{c,i}\in\mathbb{R}$: the observed value at time $t_{c,i}$. Channels may be observed at different timestamps and with different sequence lengths $N_{c}$, resulting in irregular and sparse structure when aligned over a global timeline.
Forecasting Query.
Analogous to the IMTS definition, we define a forecasting query as a collection of per-channel timepoints at which future values should be predicted:
$Q$ $\coloneqq[Q_{1},\ldots,Q_{C}]\in\mathcal{Q},\qquad\text{where}$
$Q_{c}$ $=\left(q_{c,i}\right)_{i=1}^{K_{c}}\in\textnormal{Seq}(\mathbb{R})$
Each $q_{c,i}$ specifies a future timestamp for which a value prediction is required for channel $c$. The corresponding forecasting answer is given by:
$Y$ $=[Y_{1},\ldots,Y_{C}]\in\mathcal{Y},\qquad\text{where}$
$Y_{c}$ $=\left(y_{c,i}\right)_{i=1}^{K_{c}}\in\textnormal{Seq}(\mathbb{R})$
ensuring $|Q_{c}|=|Y_{c}|=K_{c}$ for each channel $c$.
Problem 1
IMTS Forecasting Problem. Given a training dataset of $M$ triples: $\{(X^{(m)},Q^{(m)},Y^{(m)})\}_{m=1}^{M}$ drawn from an unknown distribution $\rho$ over $\mathcal{X}\times\mathcal{Q}\times\mathcal{Y}$, and a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ (e.g., mean squared error), the goal is to learn a model: $\hat{Y}:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{Y}$ that minimizes the expected loss:
$\mathcal{L}(\hat{Y};\rho):=\mathbb{E}_{(X,Q,Y)\sim\rho}\left[\ell(Y,\hat{Y}(X,Q))\right]$ (1)
5 IMTS-Mixer
Each IMTS consists of $C$ channels, where the number of observations $N_{c}$ for each channel $c$ may vary, both within a single IMTS and across different samples. This irregularity makes it challenging to directly apply standard neural architectures, particularly MLP-Mixers, which require fixed-size input representations.
To address this, we propose the Irregularly Sampled Channel Aggregation Module (ISCAM), which encodes each variable-length channel $X_{c}$ into a fixed-dimensional vector $Z_{c}\in\mathbb{R}^{D}$, enabling uniform downstream processing using MLP-Mixer blocks.
5.1 Irregularly Sampled Channel Aggregation Module
Let a single channel $X_{c}$ represent a sequence of $N_{c}$ observations. Our goal is to transform this irregular sequence into a fixed-size embedding $Z_{c}\in\mathbb{R}^{D}$. For clarity, we drop the channel subscript $c$ and denote the observations as $(t_{i},v_{i})$.
Observation-Tuple Embedding.
Each observation tuple $(t_{i},v_{i})$ is embedded using a shared MLP $f_{\textsc{ote}}:\mathbb{R}^{2}\to\mathbb{R}^{D}$, resulting in:
$H=[h_{1},\ldots,h_{N_{c}}]\in\mathbb{R}^{N_{c}\times D},\quad\text{where }h_{i}=f_{\textsc{ote}}([v_{i},t_{i}])$
Weighted Aggregation.
To compute an importance score for each observation, we apply a second shared MLP $f_{\textsc{wa}}:\mathbb{R}^{2}\to\mathbb{R}^{D}$:
$A=[a_{1},\ldots,a_{N_{c}}]\in\mathbb{R}^{N_{c}\times D},\quad\text{where }a_{i}=f_{\textsc{wa}}([v_{i},t_{i}])$
Unlike standard attention mechanisms which compute a scalar weight per observation, we assign a separate weight to each embedding feature, similar to multi-head attention, but without cross-token interactions.
To handle the variable length $N_{c}$ across samples and ensure scale invariance, we apply softmax normalization along the columns $d$:
$\text{softmax}:\mathbb{R}^{N_{c}}\to\mathbb{R}^{N_{c}},\quad\text{so that }\sum_{i=1}^{N_{c}}\text{softmax}(a_{i,d})=1$
Let $A_{:,d}=(a_{1,d},\ldots,a_{N_{c},d})$ and $H_{:,d}=(h_{1,d},\ldots,h_{N_{c},d})$ denote the $d$-th column across the $N_{c}$ elements of $A$ and $H$, respectively. The final channel embedding is then computed as:
$Z_{c}=\left[\sum_{i=1}^{N_{c}}\text{softmax}(A_{:,d})_{i}\cdot H_{:,d}\right]_{d=1}^{D}\in\mathbb{R}^{D}$ (2)
This yields a fixed-size vector per channel, where each dimension is an importance-weighted average over time.
While one can have channel-specific parameters in $f_{\textsc{ote}}$ and $f_{\textsc{wa}}$ we allow to share the parameters to reduce the number of parameters.
Channel Bias.
Since the embedding MLPs are shared across all channels, global channel-specific information may be lost. To address this, we introduce a channel bias vector $b_{c}\in\mathbb{R}^{D}$ for each channel $c$, composed of learnable parameters. The final channel encoding is computed as:
$Z_{c}^{+}=Z_{c}+b_{c}\in\mathbb{R}^{D}$ (3)
This allows the model to retain per-channel identity and compensate for missing patterns or context. In the case where a channel $c$ is missing (i.e., $X_{c}$ is empty), we define $Z_{c}:=0$ so that the output $Z_{c}^{+}=b_{c}$ contains only the bias information.
Comparison with Existing Architectures.
Several prior models such as SeFT (Horn et al. 2020), Tripletformer (Yalavarthi, Burchert, and Schmidt-Thieme 2023), and GraFITi (Yalavarthi et al. 2024a) use attention-based mechanisms over IMTS observations, where the aggregation is performed via a softmax-weighted sum of encoded tuples. These methods compute weights through the scaled dot-product of learned queries and keys, following the standard attention framework (Vaswani et al. 2017).
In contrast, ISCAM bypasses query-key interactions entirely by modeling importance weights directly through a learned MLP applied to each observation tuple. This results in a simpler and more lightweight design, while still enabling dimension-specific weighting through softmax normalization.
TTCN (Zhang et al. 2024a, b) also employs MLP-based aggregation over irregular time series. However, it introduces a more complex mechanism by using separate MLPs (called filters) for each embedding dimension $d$. Moreover, TTCN lacks a dedicated channel bias component, which limits its ability to encode global per-channel information independent of input observations.
Empirically, we show that ISCAM serves as a more effective embedding module compared to multi-head attention and TTCN in Table 5, particularly when paired with MLP-Mixer architectures for IMTS forecasting.
5.2 Mixer Blocks
Channel $\mathsf{C}\ldotsChannel$ $\mathsf{2}$Channel $\mathsf{1}$$\mathsf{Q_{C}}$$\mathsf{Q_{2}}ISCAM$$\mathsf{Q_{1}}$ISCAM
$Z_{1}$$Z_{2}$$Z_{C}$concatenate
DDCC
[…]
$\times{L}$Mixer Block
$D_{\textsc{out}}$$C$
[…]
Figure 2: Overview of IMTS-Mixer’s architecture
Similar to MLP-Mixer and TSMixer, we employ $L$ mixer blocks, each operating over the channel and feature dimensions. A single mixer block at layer $l$ updates the representation $Z^{(l-1)}\in\mathbb{R}^{C\times D}$ as follows:
$\begin{aligned} Z^{\prime{(l)}}=Z^{(l-1)}+\text{ReLU}\left(\text{Linear}_{\textsc{chan}}^{(l)}\left(\text{RMS}({Z^{(l-1)}}^{\top})\right)\right)^{\top}\\ Z^{(l)}=Z^{(l-1)}+Z^{\prime{(l)}}+\text{ReLU}\left(\text{Linear}_{\textsc{dim}}^{(l)}\left(\text{RMS}(Z^{\prime{(l)}})\right)\right)\end{aligned}$ (4)
Here, $\text{Linear}_{\textsc{chan}}^{(l)}$ operates along the channel dimension (i.e., across $C$ channels), $\text{Linear}_{\textsc{dim}}^{(l)}$ operates along the feature dimension ($D$), RMS denotes RMSNorm (Zhang and Sennrich 2019), and ReLU is the element-wise nonlinearity.
Compared to MLP-Mixer and TSMixer described in Section 3, we introduce two modifications to adapt to the irregular multivariate forecasting setting:
  • Flexible Output Dimension: We allow the output of the final mixer layer to have a configurable feature dimension $D_{\textsc{out}}$, i.e., $Z^{(L)}\in\mathbb{R}^{C\times D_{\textsc{out}}}$. This makes it possible to match the projection to varying numbers of queries per channel. All intermediate layers maintain a fixed hidden size $D$.
  • RMSNorm Instead of LayerNorm: We replace LayerNorm with RMSNorm (Zhang and Sennrich 2019), which has been shown to yield faster convergence while maintaining predictive performance.
5.3 Continuous Temporal Projection
TSMixer utilizes a linear layer to project the encoded representation of each channel to forecasts on a predefined regular grid of query steps (Chen et al. 2023). This design, however, is incompatible with IMTS, where query times are not fixed but vary continuously across instances.
To address this, IMTS-Mixer introduces a Continuous Temporal Projection (ConTP), which allows the model to project the hidden state of each channel to any given point in time. Instead of learning linear layers tied to predetermined query steps, ConTP learns a function $f_{\textsc{qu}}:\mathbb{R}\rightarrow\mathbb{R}^{D_{\textsc{out}}}$ that maps scalar query times to the weights of a linear projection dynamically. Formally, the forecast for channel $c$ at query time $q_{c,i}$ is computed as:
$\begin{aligned} \hat{y}_{i,c}&=f_{\textsc{qu}}^{c}(q_{c,i})\cdot Z_{c}^{(L)}+b^{\textsc{out}}_{c}\in\mathbb{R}\end{aligned}$ (5)
In order to enhance expressiveness, we learn a separate function $f_{\textsc{qu}}^{c}$ for each channel and implement this with a two-layer MLP.
Additionally, we learn a query independent output bias $b^{\textsc{out}}\in\mathbb{R}^{C}$.
The complete architecture of IMTS-Mixer is illustrated in Figure 2.
6 Experiments
PhysioNet MIMIC Human Activity USHCN
[…]
TSMixer 8.664 ± 0.47 4.68 ± 0.13 1.89 ± 0.01 7.99 ± 0.07 3.01 ± 0.01 3.64 ± 0.01 5.39 ± 0.04 3.33 ± 0.08
DLinear 41.86 ± 0.05 15.52 ± 0.03 4.90 ± 0.00 16.29 ± 0.05 4.03 ± 0.01 4.21 ± 0.01 6.21 ± 0.00 3.88 ± 0.02
TimesNet 16.48 ± 0.11 6.14 ± 0.03 5.88 ± 0.08 13.62 ± 0.07 3.12 ± 0.01 3.56 ± 0.02 5.58 ± 0.05 3.60 ± 0.04
PatchTST 12.00 ± 0.23 6.02 ± 0.14 3.78 ± 0.03 12.43 ± 0.10 4.29 ± 0.14 4.80 ± 0.09 5.75 ± 0.01 3.57 ± 0.02
Crossformer 6.66 ± 0.11 4.81 ± 0.11 2.65 ± 0.10 9.56 ± 0.29 4.29 ± 0.20 4.89 ± 0.17 5.25 ± 0.04 3.27 ± 0.09
GRU-D 5.59 ± 0.09 4.08 ± 0.05 1.76 ± 0.03 7.53 ± 0.09 2.94 ± 0.05 3.53 ± 0.06 5.54 ± 0.38 3.40 ± 0.28
SeFT 9.22 ± 0.18 5.40 ± 0.08 1.87 ± 0.01 7.84 ± 0.08 12.20 ± 0.17 8.43 ± 0.07 5.80 ± 0.19 3.70 ± 0.11
RainDrop 9.82 ± 0.08 5.57 ± 0.06 1.99 ± 0.03 8.27 ± 0.07 14.92 ± 0.14 9.45 ± 0.05 5.78 ± 0.22 3.67 ± 0.17
Warpformer 5.94 ± 0.35 4.21 ± 0.12 1.73 ± 0.04 7.58 ± 0.13 2.79 ± 0.04 3.39 ± 0.03 5.25 ± 0.05 3.23 ± 0.05
mTAND 6.23 ± 0.24 4.51 ± 0.17 1.85 ± 0.06 7.73 ± 0.13 3.22 ± 0.07 3.81 ± 0.07 5.33 ± 0.05 3.26 ± 0.10
Latent-ODE 6.05 ± 0.57 4.23 ± 0.26 1.89 ± 0.19 8.11 ± 0.52 3.34 ± 0.11 3.94 ± 0.12 5.62 ± 0.03 3.60 ± 0.12
Neural Flow 7.20 ± 0.07 4.67 ± 0.04 1.87 ± 0.05 8.03 ± 0.19 4.05 ± 0.13 4.46 ± 0.09 5.35 ± 0.05 3.25 ± 0.05
CRU 8.56 ± 0.26 5.16 ± 0.09 1.97 ± 0.02 7.93 ± 0.19 6.97 ± 0.78 6.30 ± 0.47 6.09 ± 0.17 3.54 ± 0.18
GraFITi 4.89 ± 0.12 3.65 ± 0.09 1.53 ± 0.02 6.68 ± 0.09 2.65 ± 0.02 3.09 ± 0.02 5.17 ± 0.11 3.22 ± 0.24
TimeCHEAT 6.18 ± 0.25 4.40 ± 0.17 1.76 ± 0.01 7.67 ± 0.06 4.42 ± 0.60 4.90 ± 0.51 5.70 ± 0.53 3.51 ± 0.31
tPatchGNN 4.98 ± 0.08 3.72 ± 0.03 1.69 ± 0.03 7.22 ± 0.09 2.66 ± 0.03 3.15 ± 0.02 5.00 ± 0.04 3.08 ± 0.04
IMTS-Mixer 4.88 ± 0.03 3.46 ± 0.01 1.61 ± 0.01 6.83 ± 0.03 2.50 ± 0.01 3.05 ± 0.01 4.91 ± 0.05 3.02 ± 0.03
Table 2: Comparison of different algorithms across four datasets based on Test MSE and MAE. We reproduced the results from GraFITi (Yalavarthi et al. 2024a) with the hyperparameters from their GitHub repository (https://github.com/yalavarthivk/GraFITi). The results for IMTS-Mixer, TSMixer and TimeCHEAT are from our own experiments. Other results are reported from tPatchGNN (Zhang et al. 2024a). Each training run was repeated 5 times with different random seeds and the standard deviation in between these trials is given after the ±. We report the best results in bold and the second-best results underlined.
We evaluate IMTS-Mixer on two benchmarks: First, we compare on the PMAU benchmark as introduced by Zhang et al. (Zhang et al. 2024a). This benchmark, contains the four datasets PhysioNet (Silva et al. 2012), MIMIC (Johnson et al. 2016) Human Activity (Vedrana Vidulin 2010) and USHCN (Menne, Williams, and Vose 2006) and we provide a more detailed description of each dataset in the appendix. Second, we use the recently published Physiome-ODE benchmark (Klötergens et al. 2024) which contains 50 semi-synthetic IMTS datasets generated from Biological ODEs. We refer to the Appendix for detailed information on hyperparameters.
6.1 PMAU
The baseline models already established in the PMAU benchmark originate from regular multivariate time series forecasting, IMTS classification and IMTS forecasting. DLinear (Zeng et al. 2023), TimesNet (Wu et al. 2022), PatchTST (Nie et al. 2022) and Crossformer (Zhang and Yan 2022) are state-of-the-art models for regular multivariate forecasting. These are applied by adding the time stamps and information about missing values as additional channels, which is shows to be an oversimplified and ineffective strategy (Zhang et al. 2024a).
GRU-D (Che et al. 2018), SeFT (Horn et al. 2020), Raindrop (Zhang et al. 2021), Warpformer (Zhang et al. 2023) and mTAND (Shukla and Marlin 2020) are all models primarily introduced for IMTS classification, which are equipped with a predictor network which inputs hidden IMTS representations and query times.
Finally, Latent-ODE (Rubanova, Chen, and Duvenaud 2019), Continuous Recurrent Units (CRU) (Schirmer et al. 2022), Neural Flow (Biloš et al. 2021), tPatchGNN (Zhang et al. 2024a) and GraFITi (Yalavarthi et al. 2024a) are models designed for IMTS Forecasting. To these baselines we add TimeCHEAT (Liu, Cao, and Chen 2025) and TS-Mixer (Chen et al. 2023) and provide the respective implementation details in Appendix B.
Table 2 reports the forecasting accuracy of IMTS-Mixer compared to the baseline models. While all models were trained on Mean Squared Error (MSE) we also report the test Mean Absolute Error (MAE). IMTS-Mixer provides the most accurate forecasts on PhysioNet, Human Activity and USHCN for both evaluation metrics MSE and MAE. For MIMIC, our model comes in second, closely behind GraFITi. We observe that GraFITi’s relative performance increases with the number of channels. While it outperforms IMTS-Mixer on MIMIC (102 channels) and is on par with our model on PhysioNet (37 channels), it is significantly outperformed by IMTS-Mixer on Human Activity (12 channels) and USHCN (5 channels).
Additionally, we observe that TimeCHEAT does not provide competitive forecasting accuracy despite being the most recent baseline (AAAI 2025). The model is likely struggling, due to its inability to account for different query time steps. This major flaw is less problematic in very short-term forecasting scenarios, as used in the paper that introduced TimeCHEAT (Liu, Cao, and Chen 2025).
6.2 Physiome-ODE
Avg. Rank # Wins %Gap-MSE
GraFITi-C 4.70 6 135
Neural Flows 7.68 0 343
CRU 4.44 2 74.4
LinODENet 3.36 7 21.1
GraFITi 3.26 2 41.6
tPatchGNN 4.78 1 79
TimeCHEAT 5.80 1 141
IMTS-Mixer 1.20 42 4.1
Table 3: We report the average rank based on Test MSE, assigning the better rank in case of ties. # Wins indicates the number of datasets where a model achieves the lowest Test MSE. %Gap-MSE refers to the mean relative difference in test MSE compared to the best model on each dataset.
In order to provide a broader evaluation we compare IMTS-Mixer on the Physiome-ODE benchmark. The 50 datasets, that are contained in Physiome-ODE were created by solving Biological ODEs for 100 steps. 80% of the resulting observations are dropped uniform at random. Each dataset consists of 2000 IMTS that each have different ODE constants and initial states. We refer to Klötergens et al. (2024) for more details. IMTS-Mixer is compared with GraFITi, LinODENet (Scholz et al. 2022), CRU, Neural Flows and GraFITi-C, a constant variant of GraFITi. Additionally, we add tPatchGNN and TimeCHEAT to the benchmark and refer to Appendix B for implementation details.
The Physiome-ODE benchmark was originally proposed to highlight the strengths of neural ODEs (Biloš et al. 2021; Schirmer et al. 2022) and in first experiments a member of that model family (LinODENet) had the best forecasting accuracy on most datasets. Now, IMTS-Mixer has the lowest test MSE on 42 of 50 datasets, establishing a new state-of-the-art on this benchmark. We list the test MSEs for all 50 datasets in Appendix C.
6.3 Efficiency Analysis
We conduct our experiments on an NVIDIA 1080Ti with 12 GB of GPU Memory. In Table 4, we compare the inference time and parameter count of recent models to that of IMTS-Mixer. Here, inference time refers to the total time it takes to answer all queries from the test set using a batch size of 32. We observe that IMTS-Mixer’s inference time is consistently shorter than the inference time of competing models. Furthermore, it uses the fewest number of parameters on three out of four datasets. The inference time and number of parameters for the Physiome-ODE datasets are listed in the Appendix C.
PhysioNet MIMIC H. Activity USHCN
GraFITi 251k / 1.9 255k / 4.0 249k / 0.6 496k / 2.2
tPATCHGNN 362k / 2.7 363k / 10 166k / 1.4 167k / 2.5
TimeCHEAT 659k / 33 755k / 72 1262k / 15 330k / 63
IMTS-Mixer 207k / 0.7 497k / 2.9 72k / 0.3 40k / 1.4
Table 4: Number of parameters (left) and inference time (right) in seconds for each dataset of the PMAU benchmark.
6.4 Evaluating Alternative Components
We conduct an ablation study in which we replace ISCAM, with Multi-Head Attention (MHA) (Vaswani et al. 2017) and Transformable Time-aware Convolution Network (TTCN) (Zhang et al. 2024b). A detailed description on how a MHA is applied in this experiment is given in Appendix D. The result is shown in Table 5. For a fair comparison, we re-tune the hyperparameters for each IMTS-Mixer modification. Our results show, that among the evaluated candidates ISCAM is the best encoder for our model.
Furthermore, we compare ConTP with using an MLP as temporal projection as implemented in tPatchGNN. Here, a sinusoidal embedding of the query times is concatenated to the hidden state of each channel and the forecast is obtained by a 2-layer MLP that inputs the enriched channel representation and outputs a scalar (Zhang et al. 2024a), On each dataset ConTP shows to be more or equally effective than the existing MLP based method.
PhysioNet MIMIC H. Activity USHCN
MSE $\times 10^{-3}$ $\times 10^{-2}$ $\times 10^{-3}$ $\times 10^{-1}$
Encoder
TTCN 5.07 ± 0.04 1.71 ± 0.05 2.54 ± 0.02 4.94 ± 0.06
MHA 4.94 ± 0.06 1.66 ± 0.01 2.52 ± 0.01 4.94 ± 0.05
ISCAM 4.88 ± 0.03 1.61 ± 0.01 2.50 ± 0.01 4.91 ± 0.05
Temporal Projection
MLP 4.91 ± 0.03 1.61 ± 0.01 2.51 ± 0.02 5.11 ± 0.03
ConTP 4.88 ± 0.03 1.61 ± 0.01 2.50 ± 0.01 4.91 ± 0.05
Table 5: Comparing Test MSE of IMTS-Mixer variants with different Encoders and Temporal Projection modules.
6.5 Effect of Mixer Blocks
We conducted an experiment to assess the effect of that the number of mixer layers has on the forecasting accuracy on the four PMAU datasets. The results shown in Figure 3 indicate that having at least one mixer block is necessary to obtain the optimal performance. However, stacking more and more mixer blocks is usually not beneficial on the evaluated datasets. Furthermore, we observe that IMTS-Mixer, without a single mixer block, is already competitive with state-of-the-art baselines, showcasing the ability of ISCAM and ConTP.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Test MSE with varying number of mixer blocks.
7 Limitations
Theoretically, IMTS-Mixer is limited by its fixed-size channel aggregation, which converts the sequence of observations from a channel into a fixed-sized embedding. For very long sequences, or when typical sequence lengths vary significantly across channels, the fixed-size representation may become a bottleneck or fail to adequately capture the diversity across channels. However, in practice, we do not find this to be an issue. Our model is able to fit all the common benchmark datasets well with varying number of channels and distribution of sequence lengths.
Another limitation is that the parameter count in the mixer blocks scales quadratically with the number of channels. We observe that IMTS-Mixer struggles with datasets that contain many channels. On the PMAU benchmark, IMTS-Mixer is outperformed by GraFITi only on the MIMIC dataset, which has the most channels and is on par with GraFITi on PhysioNet, which has the second-most channels. Furthermore, we find that MIMIC is the only dataset in which IMTS-Mixer does not have the lowest number of parameters, as shown in Table 4. However, on datasets with fewer channels (e.g., Human Activity, USHCN), our model achieves significantly higher forecasting accuracy than competing methods.
8 Conclusion and Future Work
In this work we present IMTS-Mixer, a parameter efficient an all MLP-based model for IMTS forecasting. Our architecture incorporates ISCAM, a novel method for aggregating irregularly sampled univariate time series that, despite its simplicity, outperforms existing approaches. Additionally, we generalize linear temporal projection layers to IMTS forecasting targets by introducing ConTP and yield equal or better performance than previously applied solutions.
In our experiments, we show that IMTS-Mixer establishes a new state-of-the-art forecasting accuracy on three real-world and 42 semisynthetic datasets, despite its architectural simplicity. We note that our model has difficulty handling time series with a large number of channels, as the parameter count in the fully connected layers increases significantly, which is associated with higher test error.
For future work, we plan on expanding our model to the related fields of IMTS classification and interpolation. Furthermore, we will investigate IMTS-Mixer’s applicability to probabilistic IMTS forecasting, where it could serve as an encoder of conditional normalizing flows (Yalavarthi et al. 2024b, c).
Appendix A Datasets
PhysioNet (Silva et al. 2012) and MIMIC (Johnson et al. 2016) both contain the vital signs of intensive care unit patients. To evaluate the models, we query them to forecast all the measured variables for 24 hours based on all observations from the initial 24 hours after a patient’s admission.
Human Activity (Vedrana Vidulin 2010) contains 3-dimensional positional records of four sensors attached resulting in 12 variables. The sensors are attached to individuals who perform various activities (walking, sitting among others). Models are tasked with predicting 1 second of human motion based on 3 seconds of observation.
Finally, USHCN (Menne, Williams, and Vose 2006) is created by combining daily meteorological measurements from over one thousand weather stations that are distributed over the US. While the 3 datasets mentioned above contain sparsely sampled IMTS intrinsically, USHCN originally was observed regularly and transformed into IMTS by randomly sampling observations. In each IMTS we use 2 years of observations to predict the climate conditions of a single month. For each dataset we split the samples into 60% train 20% validation and 20% test data. Additional information about the datasets are given in Table 6. We calculate the Sparsity of each dataset by dividing the number of non-missing observations ($N_{\textsc{obs}}$) by the number of given observation steps ($N_{T}$) times the number of channels ($C$): $\text{Sparsity}=\frac{N_{\textsc{obs}}}{N_{T}C}$
Previous works (De Brouwer et al. 2019; Biloš et al. 2021; Schirmer et al. 2022; Yalavarthi et al. 2024a) used the parts of these datasets, but with different preprocessing, chunking and validation protocols. We want to emphasize that therefore, the results reported in these works are incomparable.
Table 6: Summary of dataset characteristics. Obs. / Forc. Range refers to the observation range and forecasting horizon. Instances are the IMTS from training- validation- and test set combined. Channels are the number of variables observed. min/max avg. Obs. refers to the minimum/maximum number of observation per channel averaged over instances.
PhysioNet MIMIC H. Activity USHCN
Obs. / Forc. range 24h/24h 24h/24h 3s/1s 2y/1m
Sparsity 86% 97% 75% 78%
Instances 12,000 23,457 5,400 26,736
Channels 41 96 12 5
min avg. Obs. 0.07 0.0007 19.7 32.7
max avg. Obs. 31.1 3.3 23.9 35.6
Appendix B Experimental Details
IMTS-Mixer
For the PMAU benchmark, we sample 20 hyperparameter configurations and select the one with the lowest MSE on the validation split. For Physiome-ODE, we adhere to the proposed protocol and conduct random search with 10 configurations and evaluating them on the validation set of a single fold (Klötergens et al. 2024). Furthermore, we set the hidden dimension of the non-linear networks that encode observation and query time stamps to 32. To train IMTS-Mixer, we implement the schedule-free AdamW(Kingma and Ba 2017; Loshchilov and Hutter 2019; Defazio et al. 2024) optimizer with an initial learning rate of 0.01 and use early stopping with patience of 20 epochs. We use a batch-size of 32.
  • We tune the dimension of the aggregated channels $D$ from the set {64,128,256 }
  • The range for $D_{\textsc{out}}$ is {32,64,128}
  • We allow the number of mixer blocks to be {1,2,3}
  • We tune the weight decay from {1e-2, 1e-3, 1e-4}
GraFITi
We refer to the hyperparameters as given in the GitHub repository from GraFITi (Yalavarthi et al. 2024a):
https://github.com/yalavarthivk/GraFITi
  • PhysioNet: 4 layers, 4 heads, latent-dimension of 64 and a batch size of 32
  • MIMIC: 4 layers, 4 heads, latent-dimension of 64 and a batch size of 32
  • Activity: 4 layers, 4 heads, latent-dimension of 64 and a batch size of 32
  • USHCN: 2 layers, 2 heads, latent-dimension of 128 and a batch size of 32
To compare our model with GraFITi on the Physiome-ODE benchmark in terms of parameter count and inference term, we retain the hyperparameter parameters by following the search protocol as described in Klötergens et al. (2024).
TimeCHEAT
(Liu, Cao, and Chen 2025) We used the code as published in:
https://github.com/Alrash/TimeCHEAT. The batch-size is set to 32 and following the published implementation, we use a decay on plateau scheduler with Adam optimizer and an initial learning of 0.001. For TimeCHEAT we tuned the hyperparameters from the following grid:
  • Number of transformer layers: {1, 2, 3}
  • Transformer hidden dimension: {64, 128, 256 }
  • Number of encoder layers: {1, 2, 4}
  • The initial learning rate is tuned from {1e-4, 1e-3}
TSMixer
We use PyTorch implementation from TS-Mixer (Chen et al. 2023) as published here:
https://github.com/ditschuk/pytorch-tsmixer. TSMixer is applied as follows: Timestamps and a binary mask indicating whether a channel is observed at a given time step are used as additional variables, which are not forecasted. Query timestamps are input into the model as future covariates. The model is trained using the AdamW optimizer, with a weight decay of 0.0001 and a batch size of 32. We tune these following hyperparameters by random search from the these ranges:
  • The initial learning rate is tuned from {1e-4, 1e-3}
  • Dropout: {0.0, 0.1, 0.2 }
  • Number of Mixer Blocks: {1, 2, 3}
tPatchGNN
For tPatchGNN (Zhang et al. 2024a), we use the implementation as published here:
https://github.com/usail-hkust/t-PatchGNN
  • The hidden dimension is selected from $\{32,{64},128\}$
  • The time embedding dimension is taken from $\{5,{10},20\}$
  • The initial learning rate is tuned from {1e-4, 1e-3}
  • The patch size
    • is tuned from $\{{2},4,{8}\}$ for PhysioNet, MIMIC, USHCN
    • is tuned from $\{100,{300},750\}$ for Activity
Appendix C Physiome-ODE
Dataset IMTS-Mixer TimeCHEAT tPatchGNN GraFITi LinODENet CRU Neural Flows GraFITi-C
DUP01 0.954± 0.038 0.978± 0.030 0.957± 0.035 0.955± 0.037 0.964± 0.036 0.958± 0.038 1.030± 0.046 0.951± 0.036
JEL01 0.935± 0.018 0.966± 0.027 0.943± 0.016 0.942± 0.020 0.949± 0.015 0.939± 0.015 1.000± 0.013 0.935± 0.016
DOK01 0.978± 0.005 0.997± 0.004 0.987± 0.004 0.984± 0.005 0.996± 0.003 0.985± 0.005 0.998± 0.005 0.982± 0.005
INA01 1.003± 0.009 1.005± 0.010 1.004± 0.009 1.004± 0.010 1.009± 0.011 1.005± 0.010 1.008± 0.008 1.004± 0.009
WOL01 0.783± 0.026 0.804± 0.029 0.795± 0.027 0.787± 0.028 0.806± 0.027 0.814± 0.029 0.841± 0.029 0.784± 0.030
BOR01 0.708± 0.021 0.718± 0.022 0.719± 0.019 0.712± 0.022 0.719± 0.021 0.715± 0.020 0.743± 0.026 0.709± 0.022
HYN01 0.608± 0.048 0.652± 0.049 0.630± 0.045 0.625± 0.043 0.672± 0.044 0.665± 0.053 0.636± 0.045 0.619± 0.046
JEL02 0.664± 0.030 0.714± 0.027 0.702± 0.023 0.699± 0.029 0.693± 0.031 0.674± 0.028 0.779± 0.028 0.687± 0.027
DUP02 0.719± 0.046 0.748± 0.047 0.728± 0.049 0.728± 0.044 0.740± 0.042 0.722± 0.046 0.890± 0.081 0.718± 0.046
WOL02 0.640± 0.014 0.679± 0.012 0.656± 0.016 0.654± 0.014 0.663± 0.015 0.653± 0.017 0.685± 0.012 0.645± 0.016
DIF01 0.959± 0.029 0.977± 0.031 0.991± 0.027 0.985± 0.030 0.832± 0.087 0.985± 0.025 1.014± 0.025 0.982± 0.029
VAN01 0.240± 0.005 0.252± 0.007 0.247± 0.007 0.246± 0.005 0.250± 0.006 0.253± 0.005 0.250± 0.006 0.242± 0.006
DUP03 0.610± 0.050 0.683± 0.046 0.624± 0.052 0.627± 0.043 0.632± 0.044 0.622± 0.047 1.098± 0.447 0.744± 0.042
BER01 0.277± 0.013 0.311± 0.028 0.310± 0.014 0.300± 0.018 0.279± 0.020 0.280± 0.016 0.398± 0.014 0.342± 0.018
LEN01 0.547± 0.043 0.932± 0.112 0.981± 0.062 0.607± 0.055 0.387± 0.071 0.754± 0.157 1.009± 0.061 0.970± 0.063
LI01 0.177± 0.018 0.438± 0.019 0.394± 0.028 0.202± 0.013 0.084± 0.009 0.175± 0.020 0.979± 0.049 0.742± 0.010
LI02 0.360± 0.047 0.464± 0.048 0.438± 0.048 0.397± 0.058 0.434± 0.044 0.437± 0.046 0.674± 0.105 0.458± 0.056
REV01 0.603± 0.045 0.741± 0.057 0.741± 0.054 0.674± 0.055 0.597± 0.061 0.602± 0.049 0.978± 0.042 0.855± 0.050
PUR01 0.143± 0.017 0.343± 0.017 0.397± 0.082 0.153± 0.006 0.106± 0.006 0.353± 0.083 0.654± 0.102 0.476± 0.020
NYG01 0.237± 0.051 0.402± 0.061 0.357± 0.067 0.344± 0.065 0.358± 0.071 0.403± 0.092 0.442± 0.094 0.366± 0.047
PUR02 0.251± 0.022 0.474± 0.057 0.390± 0.017 0.322± 0.021 0.280± 0.028 0.293± 0.026 0.723± 0.077 0.511± 0.023
HOD01 0.383± 0.052 0.511± 0.043 0.479± 0.057 0.493± 0.046 0.441± 0.043 0.409± 0.049 0.701± 0.077 0.609± 0.056
REE01 0.031± 0.005 0.044± 0.007 0.047± 0.011 0.033± 0.007 0.045± 0.012 0.051± 0.008 0.045± 0.011 0.039± 0.012
VIL01 0.302± 0.038 0.385± 0.045 0.354± 0.028 0.344± 0.044 0.374± 0.021 0.373± 0.039 0.500± 0.060 0.378± 0.042
KAR01 0.033± 0.012 0.050± 0.013 0.048± 0.012 0.041± 0.013 0.034± 0.008 0.044± 0.012 0.069± 0.009 0.078± 0.011
SHO01 0.053± 0.007 0.085± 0.019 0.078± 0.012 0.062± 0.013 0.057± 0.006 0.095± 0.010 0.092± 0.014 0.055± 0.013
BUT01 0.180± 0.036 0.306± 0.079 0.354± 0.102 0.281± 0.071 0.254± 0.074 0.317± 0.108 0.441± 0.153 0.324± 0.091
MAL01 0.014± 0.004 0.044± 0.002 0.038± 0.011 0.020± 0.004 0.018± 0.007 0.064± 0.007 0.052± 0.002 0.054± 0.005
ASL01 0.013± 0.004 0.029± 0.003 0.031± 0.004 0.025± 0.009 0.022± 0.003 0.046± 0.014 0.066± 0.033 0.026± 0.002
BUT02 0.156± 0.022 0.272± 0.031 0.280± 0.043 0.248± 0.052 0.207± 0.056 0.282± 0.042 0.329± 0.068 0.256± 0.039
MIT01 0.003± 0.000 0.003± 0.000 0.003± 0.000 0.003± 0.000 0.003± 0.000 0.003± 0.000 0.008± 0.004 0.003± 0.000
GUP01 0.011± 0.004 0.060± 0.014 0.047± 0.016 0.041± 0.006 0.018± 0.007 0.057± 0.017 0.043± 0.017 0.035± 0.006
GUY01 0.004± 0.001 0.015± 0.004 0.005± 0.001 0.005± 0.003 0.006± 0.005 0.004± 0.001 0.024± 0.015 0.004± 0.001
PHI01 0.105± 0.022 0.308± 0.019 0.249± 0.032 0.222± 0.013 0.131± 0.014 0.133± 0.020 0.635± 0.158 0.345± 0.015
GUY02 0.007± 0.002 0.032± 0.006 0.016± 0.003 0.012± 0.009 0.010± 0.006 0.010± 0.002 0.082± 0.066 0.032± 0.015
PUL01 0.006± 0.002 0.036± 0.011 0.018± 0.006 0.008± 0.001 0.008± 0.004 0.012± 0.003 0.061± 0.060 0.024± 0.008
CAL01 0.074± 0.005 0.720± 0.357 0.212± 0.033 0.179± 0.012 0.078± 0.009 0.158± 0.008 0.867± 0.014 0.643± 0.024
WOD01 0.098± 0.007 0.372± 0.293 0.165± 0.012 0.164± 0.013 0.154± 0.016 0.113± 0.017 0.510± 0.060 0.344± 0.016
GUP02 0.397± 0.016 0.445± 0.018 0.463± 0.024 0.449± 0.027 0.469± 0.022 0.444± 0.018 0.577± 0.017 0.461± 0.025
M01 0.003± 0.000 0.013± 0.002 0.004± 0.001 0.003± 0.000 0.004± 0.001 0.005± 0.000 0.124± 0.207 0.003± 0.000
LEN02 0.031± 0.005 0.217± 0.108 0.089± 0.011 0.099± 0.021 0.039± 0.005 0.059± 0.012 0.380± 0.141 0.143± 0.022
KAR02 0.143± 0.007 0.196± 0.008 0.160± 0.009 0.151± 0.009 0.140± 0.010 0.151± 0.011 0.257± 0.016 0.252± 0.010
SHO02 0.021± 0.003 0.063± 0.015 0.070± 0.017 0.043± 0.006 0.037± 0.006 0.083± 0.015 0.109± 0.015 0.073± 0.010
MAC01 0.016± 0.002 0.046± 0.005 0.033± 0.003 0.021± 0.003 0.020± 0.003 0.065± 0.006 0.029± 0.011 0.019± 0.002
IRI01 0.027± 0.005 0.053± 0.011 0.048± 0.008 0.038± 0.017 0.037± 0.003 0.049± 0.010 0.116± 0.006 0.097± 0.008
BAG01 0.026± 0.002 0.050± 0.003 0.045± 0.003 0.029± 0.002 0.032± 0.005 0.046± 0.005 0.075± 0.012 0.109± 0.002
WOL03 0.062± 0.009 0.233± 0.053 0.142± 0.024 0.105± 0.016 0.073± 0.010 0.177± 0.016 0.479± 0.107 0.247± 0.032
WAN01 0.082± 0.007 0.170± 0.014 0.140± 0.011 0.119± 0.010 0.103± 0.012 0.125± 0.012 0.345± 0.027 0.232± 0.015
NEL01 0.006± 0.001 0.019± 0.001 0.013± 0.001 0.007± 0.000 0.010± 0.001 0.009± 0.001 0.060± 0.029 0.023± 0.006
HUA01 0.037± 0.001 0.099± 0.006 0.090± 0.012 0.063± 0.005 0.052± 0.004 0.116± 0.007 0.149± 0.015 0.115± 0.007
Table 7: Test MSE on the Physiome-ODE (Klötergens et al. 2024) benchmark. The results for the competing models are reported from the Physiome-ODE paper.
Dataset IMTS-Mixer TimeCHEAT tPatchGNN GraFITi
DUP01 173,075 1,279,382 365,594 315,329
JEL01 39,053 1,279,382 390,170 236,673
DOK01 105,219 338,203 159,610 1,255,297
INA01 139,891 337,282 159,830 15,825
WOL01 95,195 344,723 390,310 315,777
BOR01 178,201 341,702 365,614 1,253,377
HYN01 315,631 339,608 365,994 1,255,809
JEL02 17,741 645,014 365,594 627,329
DUP02 42,317 341,702 165,454 627,457
WOL02 192,745 343,695 160,414 315,649
DIF01 52,321 659,850 159,570 1,255,041
VAN01 80,459 642,011 390,330 1,254,273
DUP03 56,083 653,849 365,594 158,017
BER01 95,625 641,920 365,774 1,254,401
LEN01 42,317 646,214 365,614 315,393
LI01 65,179 343,376 365,654 315,521
LI02 48,095 343,376 365,654 1,253,633
REV01 44,699 1,282,256 390,230 1,253,633
PUR01 59,109 341,702 165,454 315,393
NYG01 139,891 343,747 165,974 1,256,705
PUR02 65,165 646,214 165,454 1,253,377
HOD01 62,141 342,531 390,210 1,253,505
REE01 39,997 655,491 365,634 315,457
VIL01 236,175 334,849 165,574 315,777
KAR01 130,321 644,156 390,450 1,255,041
SHO01 700,145 679,820 166,514 1,260,161
BUT01 61,353 342,531 390,210 1,253,505
MAL01 236,725 344,336 165,574 315,777
ASL01 143,079 651,229 390,730 1,256,833
BUT02 62,141 342,531 390,210 1,253,505
MIT01 192,745 344,082 165,534 236,993
GUP01 39,997 332,915 165,474 1,253,505
GUY01 168,461 340,631 165,434 5,249
PHI01 199,845 1,280,582 159,310 1,253,377
GUY02 168,461 653,462 390,170 236,673
PUL01 168,461 653,462 159,290 158,017
CAL01 284,951 1,282,977 390,470 1,255,169
WOD01 56,473 638,710 165,454 1,253,377
GUP02 45,205 655,491 159,330 1,253,505
M01 193,811 340,502 365,594 158,017
LEN02 176,997 638,710 159,310 1,253,377
KAR02 115,255 335,936 390,390 1,254,657
SHO02 528,233 672,187 166,514 1,260,161
MAC01 52,913 343,953 390,270 315,649
IRI01 97,937 654,482 390,590 316,673
BAG01 85,813 657,425 365,734 628,225
WOL03 275,821 658,641 365,814 1,254,657
WAN01 44,699 640,642 159,350 1,253,633
NEL01 62,141 342,789 365,634 1,253,505
HUA01 130,093 348,077 365,914 1,255,297
Table 8: Exact number of trainable parameters on each Physiome-ODE dataset.
Dataset IMTS-Mixer TimeCHEAT tPatchGNN GraFITi
DUP01 0.064 2.772 0.088 1.508
JEL01 0.061 2.704 0.125 1.419
DOK01 0.078 2.377 0.117 1.596
INA01 0.087 2.511 0.136 1.503
WOL01 0.063 2.461 0.153 1.517
BOR01 0.059 2.509 0.095 1.504
HYN01 0.092 2.526 0.154 1.601
JEL02 0.058 2.617 0.090 1.343
DUP02 0.053 2.398 0.113 1.342
WOL02 0.066 2.520 0.110 1.482
DIF01 0.071 2.623 0.107 1.592
VAN01 0.063 2.360 0.148 1.404
DUP03 0.056 2.316 0.089 1.201
BER01 0.065 2.450 0.128 1.473
LEN01 0.054 2.555 0.093 1.482
LI01 0.064 2.383 0.103 1.504
LI02 0.058 2.484 0.100 1.514
REV01 0.061 2.708 0.129 1.505
PUR01 0.060 2.530 0.122 1.531
NYG01 0.097 2.479 0.222 1.598
PUR02 0.054 2.575 0.101 1.500
HOD01 0.061 2.500 0.129 1.530
REE01 0.062 2.640 0.096 1.505
VIL01 0.070 2.375 0.164 1.508
KAR01 0.073 2.484 0.224 1.536
SHO01 0.115 2.724 0.359 1.604
BUT01 0.058 2.376 0.117 1.493
MAL01 0.066 2.445 0.146 1.484
ASL01 0.076 2.313 0.300 1.387
BUT02 0.062 2.523 0.122 1.506
MIT01 0.062 2.344 0.126 1.290
GUP01 0.056 2.284 0.118 1.390
GUY01 0.063 2.432 0.116 1.265
PHI01 0.061 2.808 0.091 1.509
GUY02 0.060 2.599 0.121 1.442
PUL01 0.060 2.576 0.089 1.356
CAL01 0.075 2.537 0.213 1.462
WOD01 0.059 2.581 0.107 1.527
GUP02 0.055 2.644 0.089 1.502
M01 0.064 2.538 0.098 1.364
LEN02 0.061 2.566 0.095 1.523
KAR02 0.065 2.479 0.242 1.503
SHO02 0.141 3.060 0.286 1.711
MAC01 0.061 2.379 0.161 1.510
IRI01 0.072 2.504 0.197 1.482
BAG01 0.057 2.368 0.113 1.260
WOL03 0.068 2.498 0.137 1.483
WAN01 0.063 2.521 0.100 1.501
NEL01 0.056 2.315 0.093 1.366
HUA01 0.075 2.514 0.137 1.568
Table 9: Inference time in seconds of each model on the datasets contained in the Physiome-ODE benchmark.
We report the test MSE of GraFITi-C (Klötergens et al. 2024), Neural Flows (Biloš et al. 2021), CRU (Schirmer et al. 2022), LinODENet(Scholz et al. 2022), GraFITi (Yalavarthi et al. 2024a), tPatchGNN (Zhang et al. 2024a), TimeCHEAT (Liu, Cao, and Chen 2025), and IMTS-Mixer on each dataset of the Physiome-ODE benchmark (Klötergens et al. 2024) in Table 7.
The number of parameters from each model for each dataset is shown in Table 8.
Additionally, we list the respective inference times in Table 9. For all models, we use an NVIDIA 1080Ti and a batch size of 32.
Appendix D Multi-Head Attention for Irregular Channels
In Section 6.4, we discuss replacing ISCAM with a single multi-head attention (Vaswani et al. 2017)(MHA)-Block. Hence, we infer the hidden state of channel $c$ with:
$Z_{c}^{\textsc{mha}}$ $=\mathcal{H}_{c}+\text{ReLU}\Bigl(\text{Linear}_{\textsc{mha}}(\mathcal{H})\Bigr),\quad\text{with}$
$\mathcal{H}_{c}$ $=\text{MHA}\Bigl(\mathcal{Q}_{c},\mathcal{K}_{c},\mathcal{V}_{c}\Bigr)$
To aggregate a channel into a fixed-sized vector, we use a singular query $\mathcal{Q}_{c}$ that is implemented with a learnable channel representation. Following previous work (Yalavarthi et al. 2024a; Liu, Cao, and Chen 2025), we use the sequence of encoded time-value tuples as keys and values ($\mathcal{K}_{c}=\mathcal{V}_{c}$). Here, these tuples are encoded by concatenating the scalar value to a sinusoidal time-embedding.
In the hyperparameter tuning of this ablation study we additionally select the number of attention heads from
{1, 2, 4, 8}.